# Modeling the dynamics of an oil drop driven by a surface acoustic wave in the underlying substrate

### v3

@@ 3.1 First order solution
The first order is associated with the propagation of acoustic waves in the fluid and hence we use the Helmholtz decomposition to separate the flow field into irrotational and divergence–free flow, i.e., $\mathbf{v}_{1}=\nabla\phi+\nabla\times\mathbf{\Psi}$, where $\phi$ is the velocity potential of the irrotational flow component, and $\mathbf{\Psi}$ represents the divergence–free flow component so that the flow vorticity is equal to $-\nabla^{2}\mathbf{\Psi}$. We ignore contributions from $\mathbf{\Psi}$, assuming $\mathbf{v}_{1}\approx\nabla\phi$  [^1] The vorticity is known to be concentrated mainly in a boundary layer flow near the solid surface of thickness $\sqrt{\mu/(\rho_{0}\omega)}$, which amounts to just a few hundred nanometers and introduces only small contributions to the bulk flow far from the solid compared with the acoustic wave described by the potential $\phi$; see Vanneste & Bühler (2011) for a general approach that further accounts for the boundary layer flow.. Here and in the following it is understood that the real part of complex equations is taken when presenting results for real quantities such as $\phi$, ${\bf v}_{1}$ or ${\bf v}_{2}$.
@@
and $k=\omega/c=2\pi/\lambda$ is the wavenumber of the sound wave in the liquid, and $\lambda$ its wavelength. The real and imaginary parts of $\kappa$ are given by
@@
The solution of Eq. (5) requires a boundary condition that captures the corresponding leading order behavior for the normal component of a Rayleigh-type-SAW at the solid surface. This condition can be derived from the solution given by Royer & Dieulesaint (1996) for the Rayleigh wave in an isotropic solid as (subscript ${\rm s}$ stands for solid),
@@
where $A$ is the displacement amplitude in the liquid. Unlike the solution in Royer & Dieulesaint (1996), the wavenumber $\kappa_{s}$ has an imaginary part, i.e.,
$\kappa_{\rm s}=k_{\rm s,r}-ik_{\rm s,i},$ (10)
due to the presence of the liquid on top of the solid that converts the simple SAW into a leaky SAW (LSAW). The consequent attenuation of the LSAW is accounted for by this imaginary part of $\kappa_{\rm s}$, which is given by (Arzt et al., 1967)
$k_{\rm s,i}=\frac{\rho_{0}}{\rho_{\rm s}}\frac{c}{c_{\rm s}^{2}}f,$ (11)
where $\rho_{\rm s}$ is the solid density, $f=\omega/2\pi$ is the SAW frequency and $c_{\rm s}$ ($=\omega/k_{\rm s,r}$) is the SAW phase velocity in the substrate. Note that the boundary condition, Eq. (9), is derived using the expansion of the nonlinear Rayleigh wave for small strains in the solid, so that the actual deformations at the solid surface as well as the inherently nonlinear relation between strain and stress in the solid contribute at higher orders of $M_{a}$, not considered here.
We comment here also on the relation between the experimentally-measured amplitude, $A_{n}$, and the theoretical one, $A$, used in Eq. (9). Although we expect $A\propto A_{n}$, it is not clear that these amplitudes are identical, as discussed in some detail by Royer & Dieulesaint (1996). A detailed analysis to determine the exact relation between $A$ and $A_{n}$ is beyond the scope of the present work, so we assume that $A=A_{n}$, a point to which we return later when discussing the comparison between experimental and theoretical results.
@@
$\rho_{0}=0.96$ g/cm3 $f=20$ MHz $k_{\rm s,r}^{-1}=0.03$ mm $k_{\rm r}^{-1}=0.01$ mm
$\mu=\mu_{b}=0.5$ Poise $\omega=125.7$ MHz $k_{\rm s,i}^{-1}=2.70$ mm $k_{\rm i}^{-1}=2.56$ mm
$\gamma=20.8$ dyn/cm $A=1.04-10$ nm $\lambda_{\rm s}=194$ $\mu$m $\lambda=67.5$ $\mu$m
$c=1350$ m/s $A_{n}=0-1.69$ nm $\delta=0.91$ $\mu$m
@@
This approximate value of $k_{\rm i}$ is in agreement with that defined in Eq. (5) of Brunet et al. (2010c). Note that $\beta\ll 1$ for a range of fluids; e.g., for water ($\nu=1$ cSt) and oil ($\nu=50$ cSt) we obtain $\beta=1.67\times 10^{-4}$ and $\beta=8.38\times 10^{-3}$, respectively, assuming $\nu_{b}=\nu$ in both cases. Therefore, we can consider the ratio $k_{\rm i}/k_{\rm r}$ as a small parameter, as discussed further in Appendix B.
@@
$v_{1,x}$ $=$ $\frac{\partial\phi}{\partial x}=\frac{i\kappa_{s}}{\sqrt{\kappa_{s}^{2}-\kappa^{2}}}A\omega\exp({i\omega t})\exp{\left(-i\kappa_{s}x\right)}\exp{\left(-z\sqrt{\kappa_{s}^{2}-\kappa^{2}}\right)},$
$v_{1,z}$ $=$ $\frac{\partial\phi}{\partial z}=A\omega\exp({i\omega t})\exp{\left(-i\kappa_{s}x\right)}\exp{\left(-z\sqrt{\kappa_{s}^{2}-\kappa^{2}}\right)}.$ (14)
@@ 3.2 Second order solution
Having obtained $\mathbf{v}_{1}$, we next seek the quasi–steady flow field at times long compared to the acoustic period (Stokes, 1847; Rayleigh, 1884; Schlichting, 1932). Since we excite the fluid using a single-frequency SAW, it is sufficient to time–average the system of equations over the period of the SAW (Vanneste & Bühler, 2011) using the operator $\equiv\omega/(2\pi)\int_{t=0}^{2\pi/\omega}\cdot~{}dt$. The leading order velocity contribution that supports a steady flow component (and hence does not vanish) is $\mathbf{v}_{2}\right>$, appearing at order $M_{a}^{2}$ in the expansions specified by Eq. (3). The corresponding conservation of momentum and mass equations are then given by
[…]
@@
$\mathbf{F}_{s}=-\langle\rho_{0}\left(\mathbf{v}_{1}\cdot\nabla\right)\mathbf{v}_{1}+\mathbf{v}_{1}\nabla\cdot\left(\rho_{0}\mathbf{v}_{1}\right)\rangle,$ (17)
represents the acoustic force on the induced steady streaming flow. Moreover, considering the component of the continuity equation proportional to $M_{a}$, $\partial\rho_{1}/\partial t+\rho_{0}\nabla\cdot{\bf v}_{1}=0$, and the solution for ${\bf v}_{1}$ specified by Eq. (3.1), we see that the product $\rho_{1}{\bf v}_{1}$ leads to an odd function over the SAW period, which results in $\rho_{1}{\bf v}_{1}\right>=0$. Hence, Eqs. (15) and (16) may be simplified to
[…]
${\bf v}_{2}\right>=0.$ (19)
In Appendix A we use the real parts of the first order velocity field, $\mathbf{v_{1}}$, to reveal the form of the SAW forcing as given above, and in Appendix B we discuss a simplified version as well as the inviscid limit.
@@ 3.3 Long-wave approximation
We carry out a standard long–wave expansion of Eqs. (18) and (19), focusing on the simplified two-dimensional geometry that assumes translational invariance in the transverse $y$-direction ($\partial/\partial y=0$). The discussion of three–dimensional effects will be presented elsewhere. Here, we obtain the pressure, $p\equiv\langle p_{2}\rangle$, by substituting the $z$-component of ${\mathbf{F}}_{s}$, given in Eqs. (47) and (48), in the $z$ component of Eq. (18). We then integrate the resulting equation over the film thickness with respect to $z$, using the Laplace pressure boundary condition $p(x,h(x,t))=-\gamma h^{\prime\prime}$ (simplifying the curvature of the free surface in the spirit of the long–wave approximation, and using $\prime$ to denote $\partial/\partial x$), to find
@@
$\psi(x,z)=e^{-2(k_{s,i}x+K_{z}z)},\qquad P_{0}=\rho_{0}A^{2}\omega^{2}.$ (21)
The last term on the right hand side of Eq. (20) is the contribution of the bulk acoustic forcing to the local pressure in the film. To obtain the velocity component in the $x$–direction at this order, $v_{2,x}$, we substitute both the $x$–derivative of the pressure in Eq. (20), and $F_{s,x}$ as given in Eq. (47), into the $x$–component of Eq. (18). We then integrate twice with respect to $z$, applying a zero shear-stress boundary condition at the free surface $(\partial v_{2,x}/\partial z=0$ at $z=h$) and a no-slip condition at the substrate $(v_{2,x}=0$ at $z=0)$, which yields
$v_{2,x}=$ $-\frac{1}{2\mu}z(z-2h)\left[\gamma h^{\prime\prime\prime}-\rho_{0}gh^{\prime}+C_{z}P_{0}\psi(x,h)\left(\frac{k_{\rm s,i}}{K_{z}}+h^{\prime}\right)\right]$
$\quad-\mathcal{C}\frac{P_{0}}{4\mu K_{z}^{3}}e^{-2k_{\rm s,i}x}\left(e^{-2K_{z}z}+2K_{z}e^{-2K_{z}h}z-1\right),$ (22)
@@
with parameter $C_{x}$ defined in (49). Note that the last term in Eq. (22) leads to a velocity profile $v_{2,x}$ that is modified from the usual parabolic shape in the neighborhood of $z=0$. Interestingly, there is a critical value of $\mathcal{C}$, namely $\mathcal{C}_{\rm crit}<0$, such that for $\mathcal{C}<\mathcal{C}_{\rm crit}$, the velocity profile includes negative values of $v_{2,x}$ in the region near the substrate. We have not, however, observed such flow inversion with the present choice of parameters.
From the velocity profile above, Eq. (22), we obtain the film-averaged velocity and the flux as
[…]
@@
For our simulations, we nondimensionalize the problem using an arbitrary length scale, $\ell$, a timescale $t_{c}=3\mu\ell/\gamma$ such that leading order terms balance in Eq. (25), and a pressure scale $p_{c}=\gamma/\ell$ based on the capillary contribution in Eq. (27), so that we have
$(x,h)=\ell(\tilde{x},\tilde{h}),\quad t=\frac{3\mu\ell}{\gamma}\tilde{t},\quad{\mathcal{P}}=\frac{\gamma}{\ell}\tilde{{\mathcal{P}}},\quad(k_{\rm s,i},K_{z},C_{x},C_{z})=\ell^{-1}(\tilde{k}_{\rm s,i},\tilde{K}_{z},\tilde{C}_{x},\tilde{C}_{z}),\quad{\cal C}=\ell^{-2}\tilde{\cal C}.$ (28)
@@
$\frac{\partial\tilde{h}}{\partial\tilde{t}}+\frac{\partial}{\partial\tilde{x}}\bigg{[}-\tilde{h}^{3}\frac{\partial\tilde{{\mathcal{P}}}}{\partial\tilde{x}}-\tilde{\mathcal{C}}\frac{3\mathcal{S}}{8\tilde{K}_{z}^{4}}\psi(x,h)\Big{(}2\tilde{K}_{z}^{2}\tilde{h}^{2}-1+e^{2\tilde{K}_{z}\tilde{h}}(1-2\tilde{K}_{z}\tilde{h})\Big{)}\bigg{]}=0,$ (29)
where the dimensionless effective pressure is
$\tilde{\mathcal{P}}=-\tilde{h}^{\prime\prime}+{\rm Bo}\tilde{h}+\frac{\mathcal{S}\tilde{C}_{z}}{2\tilde{K}_{z}}\psi(\tilde{x},\tilde{h}),$ (30)

### v4

@@ 3.1 First order solution
The first order is associated with the propagation of acoustic waves in the fluid and hence we use the Helmholtz decomposition to separate the flow field into irrotational and divergence–free flow, i.e., $\mathbf{v}_{1}=\nabla\phi+\nabla\times\mathbf{\Psi}$, where $\phi$ is the velocity potential of the irrotational flow component, and $\mathbf{\Psi}$ represents the divergence–free flow component, so that the flow vorticity is equal to $\nabla^{2}\mathbf{\Psi}$. We ignore contributions from $\mathbf{\Psi}$, assuming $\mathbf{v}_{1}\approx\nabla\phi$  [^1] The vorticity is known to be concentrated mainly in a boundary layer flow near the solid surface of thickness $\sqrt{\mu/(\rho_{0}\omega)}$, which amounts to just a few hundred nanometers and introduces only small contributions to the bulk flow far from the solid compared with the acoustic wave described by the potential $\phi$; see Vanneste & Bühler (2011) for a general approach that further accounts for the boundary layer flow.. Here and in the following, it is understood that the real part of complex equations is taken when presenting results for real quantities such as $\phi$, ${\bf v}_{1}$ or ${\bf v}_{2}$.
@@
and $k=\omega/c=2\pi/\lambda$ is the wavenumber of the sound wave in the liquid, with wavelength $\lambda$. The real and imaginary parts of $\kappa$ are given by
@@
The solution of Eq. (5) requires a boundary condition that captures the corresponding leading order behavior for the normal component of a Rayleigh-type-SAW at the solid surface. This condition can be derived from the solution given by Royer & Dieulesaint (1996) for the Rayleigh wave in an isotropic solid as
@@
where subscript $\rm s$} stands for solid, and $A$ is the displacement amplitude in the liquid. Unlike the solution in Royer & Dieulesaint (1996), the wavenumber $\kappa_{\rm s}$ has an imaginary part, i.e.,
$\kappa_{\rm s}=k_{\rm s,{\rm r}}-ik_{\rm s,{\rm i}},$ (10)
due to the presence of the liquid on top of the solid that converts the simple SAW into a leaky SAW (LSAW). The consequent attenuation of the LSAW is accounted for by the imaginary part of $\kappa_{\rm s}$, which is given by (Arzt et al., 1967)
$k_{\rm s,{\rm i}}=\frac{\rho_{0}}{\rho_{\rm s}}\frac{c}{c_{\rm s}^{2}}f,$ (11)
where $\rho_{\rm s}$ is the solid density, $f=\omega/2\pi$ is the SAW frequency and $c_{\rm s}$ ($\omega/k_{\rm s,{\rm r}}$) is the SAW phase velocity in the substrate. Note that the boundary condition, Eq. (9), is derived using the expansion of the nonlinear Rayleigh wave for small strains in the solid, so that the actual deformations at the solid surface as well as the inherently nonlinear relation between strain and stress in the solid, contribute at higher orders of $M_{a}$, not considered here.
We comment here also on the relation between the experimentally-measured amplitude, $A_{\rm n}$, and the theoretical one, $A$, used in Eq. (9). Although we expect $A\propto A_{\rm n}$, it is not clear that these amplitudes are identical, as discussed in some detail by Royer & Dieulesaint (1996). A detailed analysis to determine the exact relation between $A$ and $A_{\rm n}$ is beyond the scope of the present work, so we assume that $A=A_{\rm n}$, a point to which we return later when discussing the comparison between experimental and theoretical results.
@@
$\rho_{0}=0.96$ g/cm3 $f=20$ MHz $k_{\rm s,{\rm r}}^{-1}=0.03$ $m$ $k_{\rm r}^{-1}=0.01$ $m$
$\nu=\nu^{\prime}=50$ cSt $\omega=125.7$ MHz $k_{\rm s,{\rm i}}^{-1}=2.70$ $m$ $k_{\rm i}^{-1}=2.56$ $m$
[…]
@@
This approximate value of $k_{\rm i}$ is in agreement with that defined in Eq. (5) of Brunet et al. (2010). Note that $\beta\ll 1$ for a range of fluids; e.g., for water ($\nu=1$ cSt) and oil ($\nu=50$ cSt) we obtain $\beta=1.67\times 10^{-4}$ and $\beta=8.38\times 10^{-3}$, respectively, assuming $\nu_{b}=\nu$ in both cases. Therefore, we can consider the ratio $k_{\rm i}/k_{\rm r}$ as a small parameter, as discussed further in Appendix B.
@@
$v_{1,x}$ $=$ $\frac{\partial\phi}{\partial x}=\frac{i\kappa_{\rm s}}{\sqrt{\kappa_{\rm s}^{2}-\kappa^{2}}}A\omega\exp({i\omega t})\exp{\left(-i\kappa_{\rm s}x\right)}\exp{\left(-z\sqrt{\kappa_{\rm s}^{2}-\kappa^{2}}\right)},$
$v_{1,z}$ $=$ $\frac{\partial\phi}{\partial z}=A\omega\exp({i\omega t})\exp{\left(-i\kappa_{\rm s}x\right)}\exp{\left(-z\sqrt{\kappa_{\rm s}^{2}-\kappa^{2}}\right)}.$ (14)
@@ 3.2 Second order solution
Having obtained $\mathbf{v}_{1}$, we next seek the quasi–steady flow field at times long compared to the acoustic period (Stokes, 1847; Rayleigh, 1884; Schlichting, 1932). Since we excite the fluid using a single-frequency SAW, it is sufficient to time–average the system of equations over the period of the SAW (Vanneste & Bühler, 2011) using the operator $\left<\cdot\right>\equiv\omega/(2\pi)\int_{t=0}^{2\pi/\omega}\cdot~dt$. The leading order velocity contribution that supports a steady flow component (and hence does not vanish) is $\left<\mathbf{v}_{2}\right$>, appearing at order $M_{a}^{2}$ in the expansions specified by Eq. (3). The corresponding conservation of momentum and mass equations are then given by
$-\nabla\left<p_{2}\right>+{\mathbf{F}}_{\rm s}-\rho_{0}g\mathbf{e}_{z}+\mu\nabla^{2}\left<\mathbf{v}_{2}\right>+\left(\mu_{b}+\frac{\mu}{3}\right)\nabla\left(\nabla\cdot\left<\mathbf{v}_{2}\right>\right)=0,$ (15)
$\rho_{0}\nabla\cdot\left<{\bf v}_{2}\right>+\nabla\cdot\left<\rho_{1}{\bf v}_{1}\right>=0,$ (16)
@@
$\mathbf{F}_{\rm s}=-\langle\rho_{0}\left(\mathbf{v}_{1}\cdot\nabla\right)\mathbf{v}_{1}+\mathbf{v}_{1}\nabla\cdot\left(\rho_{0}\mathbf{v}_{1}\right)\rangle,$ (17)
represents the acoustic force on the induced steady streaming flow. Moreover, considering the component of the continuity equation proportional to $M_{a}$, $\partial\rho_{1}/\partial t+\rho_{0}\nabla\cdot{\bf v}_{1}=0$, and the solution for ${\bf v}_{1}$ specified by Eq. (3.1), we see that the product $\rho_{1}{\bf v}_{1}$ leads to an odd function over the SAW period, which results in $\left<\rho_{1}{\bf v}_{1}\right>=0$. Hence, Eqs. (15) and (16) may be simplified to
$-\nabla\left<p_{2}\right>+{\mathbf{F}}_{\rm s}-\rho_{0}g\mathbf{e}_{z}+\mu\nabla^{2}\left<\mathbf{v}_{2}\right>=0,$ (18)
$\nabla\cdot\left<{\bf v}_{2}\right>=0,$ (19)
with ${\mathbf{F}}_{\rm s}$ as specified in (17). In Appendix A we use the real part of the first order velocity field, $\mathbf{v_{1}}$, to reveal the form of the SAW forcing as given in Eq. (17) above, and in Appendix B we discuss a simplified version as well as the inviscid limit. The main points we emphasize are that the acoustic streaming force, given by Eq. (17), is nonconservative in contrast to the result in Shiokawa et al. (1989), and that its prefactor is significantly dependent on the presence of viscous dissipation. The discussion is made explicit in Appendix A where the real component is shown along with the exact form of the coefficients therein.
@@ 3.3 Long-wave approximation
We carry out a standard long–wave expansion of Eqs. (18) and (19), focusing on the simplified two-dimensional geometry that assumes translational invariance in the transverse $y$-direction ($\partial/\partial y=0$). The discussion of three–dimensional effects will be presented elsewhere. Here, we obtain the pressure, $p\equiv\langle p_{2}\rangle$, by substituting the $z$-component of ${\mathbf{F}}_{\rm s}$, given in Eqs. (48) and (49), in the $z$ component of Eq. (18). We then integrate the resulting equation over the film thickness with respect to $z$, using the Laplace pressure boundary condition $p(x,h(x,t))=-\gamma h^{\prime\prime}$ (simplifying the curvature of the free surface in the spirit of the long–wave approximation, and using $\prime$ to denote $\partial/\partial x$), to find
@@
$\psi(x,z)=e^{-2(k_{{\rm s},{\rm i}}x+K_{z}z)},\qquad P_{0}=\rho_{0}A^{2}\omega^{2}.$ (21)
The last term on the right hand side of Eq. (20) is the contribution of the bulk acoustic forcing to the local pressure in the film. To obtain the velocity component in the xx–direction at this order, $v_{2,x}$, we substitute both the xx–derivative of the pressure in Eq. (20), and $F_{{\rm s},x}$ as given in Eq. (48), into the xx–component of Eq. (18). We then integrate twice with respect to $z$, applying a zero shear-stress boundary condition at the free surface $(\partial v_{2,x}/\partial z=0$ at $z=h$) and a no-slip condition at the substrate $(v_{2,x}=0$ at $z=0)$, which yields
$v_{2,x}=$ $-\frac{1}{2\mu}z(z-2h)\left[\gamma h^{\prime\prime\prime}-\rho_{0}gh^{\prime}+C_{z}P_{0}\psi(x,h)\left(\frac{k_{\rm s,{\rm i}}}{K_{z}}+h^{\prime}\right)\right]$
$\quad-\mathcal{C}\frac{P_{0}}{4\mu K_{z}^{3}}e^{-2k_{\rm s,{\rm i}}x}\left(e^{-2K_{z}z}+2K_{z}e^{-2K_{z}h}z-1\right),$ (22)
@@
with parameter $C_{x}$ defined in (50). Note that the last term in Eq. (22) leads to a velocity profile $v_{2,x}$ that is modified from the usual parabolic shape in the neighborhood of $z=0$. Interestingly, there is a critical value of $\mathcal{C}$, namely $\mathcal{C}_{\rm crit}<0$, such that for $\mathcal{C}<\mathcal{C}_{\rm crit}$, the velocity profile includes negative values of $v_{2,x}$ in the region near the substrate. We have not, however, observed such flow inversion with the present choice of parameters relevant to our experiments.
From the velocity profile above, Eq. (22), we obtain the film-averaged velocity and flux as
$u=\frac{1}{h}\int\limits_{0}^{h}v_{2,x}\ dz,\qquad Q=uh.$ (24)
@@
For our simulations, we nondimensionalize the problem using an arbitrary length scale, $\ell$, a timescale $t_{c}=3\mu\ell/\gamma$ such that leading order terms balance in Eq. (25), and a pressure scale $p_{\rm c}=\gamma/\ell$ based on the capillary contribution in Eq. (27), so that we have
$(x,h)=\ell(\tilde{x},\tilde{h}),\quad t=\frac{3\mu\ell}{\gamma}\tilde{t},\quad{\mathcal{P}}=\frac{\gamma}{\ell}\tilde{{\mathcal{P}}},\quad(k_{\rm s,{\rm i}},K_{z},C_{x},C_{z})=\ell^{-1}(\tilde{k}_{\rm s,{\rm i}},\tilde{K}_{z},\tilde{C}_{x},\tilde{C}_{z}),\quad{\cal C}=\ell^{-2}\tilde{\cal C}$. (28)
@@
$\frac{\partial\tilde{h}}{\partial\tilde{t}}+\frac{\partial}{\partial\tilde{x}}\bigg[-\tilde{h}^{3}\frac{\partial\tilde{{\mathcal{P}}}}{\partial\tilde{x}}-\tilde{\mathcal{C}}\frac{3\mathcal{S}}{8\tilde{K}_{z}^{4}}\psi(\tilde{x},\tilde{h})\Big(2\tilde{K}_{z}^{2}\tilde{h}^{2}-1+e^{2\tilde{K}_{z}\tilde{h}}(1-2\tilde{K}_{z}\tilde{h})\Big)\bigg]=0,$ (29)
[…]
@@
We pause here to provide a justification for neglecting the effects of acoustic radiation pressure in the present model. The experiments modeled here involve a SAW propagating in the solid with attenuation length of 11-22 mm, while the drops considered are 66-77 mm in lateral length. When comparing the relative strengths of acoustic streaming and acoustic radiation pressure, an important distinction is that streaming results from the SAW in the solid, while acoustic radiation pressure results from the leaky SAW in the liquid that reflects off the drop’s free surface. The attenuation lengths of both effects are proportional to the relative phase velocities of the acoustic fields in the solid and fluid, respectively. Since the speed of sound in the liquid is 22-33 times smaller than the speed of sound in the solid, the radiation pressure attenuates twice as fast as the SAW in the solid. Hence, the radiation pressure affects the drop rear only along a distance of approximately 0.50.5-11 mm. Because the present focus is on describing the front dynamics of driven drops, we neglect this contribution. However, we note that in cases where the attenuation length of the SAW is comparable to the lateral length of the drop, acoustic radiation pressure is expected to have an effect on the drop geometry and dynamics comparable to that of acoustic streaming, and hence should be included (see, e.g., Marcos et al. (2025)). To quantify the regime in which radiation pressure becomes important, some insight can be reached by considering the nondimensional ratio
$R=\frac{1}{2r_{\rm d}k_{\rm i}},$ (32)
that is, the attenuation length of the SAW in the solid divided by the lateral length of the drop. When $R=\mathcal{O}(1)$, acoustic radiation pressure should be considered to properly model the drop front, but when $R\ll 1$ (e.g., $R=1/5$ as in our case), it can be neglected when modeling the drop front dynamics.